OpenStudy (anonymous):

What is the length of line DE? A 15 b 10 c 6 d 9

OpenStudy (anonymous):



OpenStudy (anonymous):

You can use proportions for this.

OpenStudy (anonymous):

I tried. I still don't understand.

OpenStudy (anonymous):

\[\frac{ 12 }{ 18 } = \frac{ x }{ 15 }\]

OpenStudy (anonymous):

Cross multiply. 12 * 15 = 18x

OpenStudy (anonymous):

so it would be 15?

OpenStudy (anonymous):

180 = 18x x = 10

OpenStudy (anonymous):

wait my teacher helped me with this also and I got 10 but she said the answer was incorrect.

OpenStudy (anonymous):

So the answer would be B) 10

OpenStudy (anonymous):

No I just got the wrong answer

OpenStudy (anonymous):

B is correct